density of a noble gas is 2.71g/l at 3.00atm and 0.00C. identify the gas
Answer #1
      Given, density of noblegas = 2.71 g/L
     
                     Pressure P = 3.00 atm
                         Temperature = T = 0.0 oC = 273 K
             Theideal gas equation is   PV = nRT
                                                          = (m /M)RT
                                                    P*M  = (m/V)RT    ( m/V = density of gas)
                                              ∴   P*M  =d RT ( M = molar mass of gas)
                            ∴ 3.0 atm * M = 2.71 g/L * 0.0821L.atm/mol.K *273K
                             ∴ M = 20.24 g/mol
                 The noble gas with molar mass 20.24 g/mol is Neon.